OpenStudy (anonymous):

Monroe is graphing the equation . He has placed a point on the origin, and uses the slope to find two more points. Which points does Monroe plot on his graph? Choose exactly two answers that are correct. A. (−9, 4) B. (−4, 9) C. (4, −9) D. (9, −4)

OpenStudy (anonymous):

Will fan :]

pooja195 (pooja195):

lol i was thinking about your problem :) and i think you left out the equation

OpenStudy (anonymous):

sorry !

pooja195 (pooja195):

slope = -4/9 slope = rise / run = -4/9 you can choose rise = -4 and run = +9 or rise = +4 and run = -9.

OpenStudy (anonymous):

so 14, 9 ?

OpenStudy (anonymous):

-4 i mean

pooja195 (pooja195):

no positive so (4,-9)

pooja195 (pooja195):

rise is along the y-axis and run is along the x-axis. start from the origin: case 1) rise = -4 and run = +9 x = 9, y = -4. Point is (9, -4) case 2) rise = +4 and run = -9 x = -9, y = 4. Point is (-9, 4)

pooja195 (pooja195):

so A and D

OpenStudy (anonymous):

ohhh (: and the second one ? -9,4 ?

OpenStudy (anonymous):

ohh okay thanks soo much

pooja195 (pooja195):

:) no problem message me anytime :) u need help

OpenStudy (anonymous):

thanks!! @pooja195 i do have a few more but ill try them

pooja195 (pooja195):

ok well message me or post them if you don't get em i'll be glad to help :)

pooja195 (pooja195):

1st one

OpenStudy (anonymous):

@pooja195 thanks !!

OpenStudy (anonymous):

What is the slope of the hypotenuses of the triangles in simplest form? http://static.k12.com/calms_media/media/1503500_1504000/1503532/1/2f86043497482838775aa2b5f0bd5f65a773fbdd/MS_IMC-140523-130901.jpg Can you help with another please (: @pooja195

pooja195 (pooja195):

repost it as a new question im not sure of it :) sorry!